What is the mass fraction and volume fraction of an aqueous solution having 0.7304 Mol Fraction, 0.6268 Specific Gravity of solute and 42.00 Molar Weight of solute?


1
Expert's answer
2021-10-07T03:44:41-0400

Molar mass of solvent = 42/ 0.6268


= 66.8789


Total mass= 66.8789 + 42

= 108.8789


Mass fraction= 42/108.8789

= 0.3857


Volume fraction = 0.3857/ 0.7304


=0.528